# Hierarchical Adaptive networks with Task vectors for Test-Time Adaptation

### v1

@@ Acknowledgments
This work is supported by the Munich Center for Machine Learning and the DAAD programme Konrad Zuse Schools of Excellence in Artificial Intelligence, both sponsored by the Federal Ministry of Research, Technology and Space. DML and JAS received funding from HELMHOLTZ IMAGING, a platform of the Helmholtz Information and Data Science Incubator.
Supplementary Material
@@ A.2 Modification of Source Training
The test-time adaptation landscape comprises two primary approaches: (i) Fully test-time adaptation methods and (ii) Methods that modify source training (often termed training preparation [88]). Fully test-time adaptation methods [83, 48] leave source training unaltered, and in line with this, our three proposed innovations are applied exclusively at test time. Common methods, however, often rely on single-dimensional architectures that fail to capture multi-scale learning across the model. Consequently, we enhance common Test-time adaptation methods to tackle diverse distribution shifts by incorporating hierarchical linear layers during source training. Similarly, other Test-time adaptation approaches [90, 91, 57, 22, 66] also modify source training to learn strategies for adaptation.
@@ A.8 Motivation for not using efficient and fixed-feature versions from Matryoshka representation learning
Algorithms We provide the details about source model training in Algorithm 2. The detailed test-time algorithm is in the main paper.
@@ B.2 Results across common test-time adaptation setting
Table 6: Comparisons on common test-time adaptation setting. for CIFAR10-C and CIFAR100-C. Integrating Hi-Vec improves the common methods and achieves the best results (bold).
@@ B.3 Inference time
During source training, Hi-Vec introduces a small parameter overhead of approximately 5% compared to standard implementations. At test-time, Hi-Vec needs to compute multiple gradient norms for layer selection. However, to adapt to target data, it doesn’t introduce any additional backward passes or parameters needed for backpropagation. Specifically, Hi-Vec uses only one of the hierarchical linear layers for backpropagation and during inference, ensuring no additional parameters are introduced. In Table 7, we compare the inference times for common test-time adaptation methods with and without Hi-Vec integration based on a single Nvidia A100 GPU. Hi-Vec introduces a modest increase in compute time for Tent and SAR, with Tent + Hi-Vec taking 1m 25s compared to 31s for Tent and SAR + Hi-Vec taking 1m 42s compared to 50s for SAR. For STAMP, Hi-Vec results in a negligible change. Overall, Hi-Vec achieves robust adaptation with minimal computational overhead, making it useful for test-time adaptation.
@@ C.1 Additional datasets information
We train source models respectively on the source datasets as per the training and evaluation procedure as in  [83, 48, 97, 27]. Akin to these common methods, the source datasets to train the respective source models include CIFAR-10, CIFAR-100, Waterbirds, and ImageNet as datasets. We utilize ResNet-18 and ResNet-50 with batch norm as the encoders. Our experimental setup involves multiple datasets at test time to evaluate the proposed method under two distinct types of diverse distribution shifts at test time. First, we consider outlier-aware scenarios where each test batch contains samples from Cifar-10-c [35], Cifar-100-c [35], or ImageNet-C [35] datasets combined with a significant proportion of outlier datasets, including LSUN-C [97], SVHN-C [97], TinyImageNet-C [97], Places-365-C [97], and Textures-C [97]. For the in-distribution data, Cifar-10-c [35] and Cifar-100-c [35] include 15 corruption types (e.g., Gaussian noise, motion blur, fog) applied at five severity levels, with Cifar-10-c [35] containing 10,000 images per corruption and Cifar-100-c [35] containing 10,000 images per class. ImageNet-C [35] serves as a large-scale benchmark with similar corruption types across 1,000 categories. We follow [97] to generate the outlier datasets. The outlier datasets are derived by applying the same 15 corruption types at the highest severity level of 5 to datasets such as LSUN, SVHN, TinyImageNet, Places-365, and Textures. In this setup, a single test batch contains both in-distribution samples from the corrupted CIFAR or ImageNet datasets and out-of-distribution samples from these corrupted outlier datasets. Next, we evaluate robustness to spurious correlations using datasets specifically designed to introduce misleading or non-causal correlations. The Waterbirds [73] dataset consists of 11,788 images where spurious correlations arise due to the background (e.g., water or land) being associated with specific bird species. Similarly, ColoredMNIST [73] modifies the MNIST dataset by introducing color as a spurious feature correlated with digit labels, consisting of 70,000 images across 10-digit classes.
@@ Appendix D Additional Related work
In addition to the aforementioned downstream tasks, test-time adaptation has also been used in vision language models [6, 103, 79, 1], continual learning [38, 58, 96, 3], classification that consider practical scenarios [28, 14, 17, 43, 81, 50, 21, 3, 4]. Moreover, recently [42], proposed a framework that uses linear mode connectivity for adapted models during inference, with standard ResNet models.

### v2

@@ Acknowledgments
This work is supported by DAAD programme Konrad Zuse Schools of Excellence in Artificial Intelligence and the Munich Center for Machine Learning, both sponsored by the Federal Ministry of Research, Technology and Space. DML and JAS received funding from HELMHOLTZ IMAGING, a platform of the Helmholtz Information and Data Science Incubator. MH is in part supported by the Munich School of Data Science (MuDS) and the European Laboratory for Learning and Intelligent Systems (ELLIS) PhD program. LD is supported but the German Federal Ministry of Research, Technology and Space (DECIPHER-M, 01KD2420G).
@@ A.2 Modification of Source Training
The test-time adaptation landscape comprises two primary approaches: (i) Fully test-time adaptation methods and (ii) Methods that modify source training (often termed training preparation [87]). Fully test-time adaptation methods [82, 47] leave source training unaltered, and in line with this, our three proposed innovations are applied exclusively at test time. Common methods, however, often rely on single-dimensional architectures that fail to capture multi-scale learning across the model. Consequently, we enhance common TTA methods to tackle diverse distribution shifts by incorporating hierarchical linear layers during source training. Similarly, other TTA approaches [89, 90, 56, 21, 65] also modify source training to develop strategic priors for adaptation.
@@ A.8 Motivation for not using efficient and fixed-feature versions from Matryoshka representation learning
A.9 Additional Notations for Merging Layer Weights of Small and Large Linear Layers
Let $W_{\phi}$ denote the weight matrix of layer $\phi$ with dimensions $(m\times n)$, and let $W_{\phi^{*}}$ denote the weight matrix of the dynamically selected layer $\phi^{*}$ with dimensions $(k\times l)$. These matrices can be written explicitly as,
$W_{\phi}=\left(w_{\phi,ij}\right)_{i=1,\dots,m;\;j=1,\dots,n},$
$W_{\phi^{*}}=\left(w_{\phi^{*},ij}\right)_{i=1,\dots,k;\;j=1,\dots,l}.$
For dimension-aware merging, we define the shared dimensions
$g=\max(m,k),\quad h=\min(n,l),$
where $\max$ selects the larger dimension and min selects the smallest dimension in respective directions. Then, the weights of $W_{\phi}$ are projected into the common shape $(g\times h)$, aligned with the indexing of $W_{\phi^{*}}$:
$W_{\phi}[\phi^{*}]:=\left(w_{\phi,ij}\right)_{i=1,\dots,g;\;j=1,\dots,h}.$
A.10 Extending to Parameter-free baselines
We also extend Hi-Vec to parameter-free baselines such as Laplacian Adjusted Maximum-likelihood Estimation (LAME) [7], which adjusts the model’s output for target batches. Specifically, we integrate Hi-Vec’s dynamic layer selection to perform targeted inference for each batch instead of representations from a static linear layer. Next, leveraging hierarchical layer agreement to decide whether the model’s output adjustment with LAME [7] is necessary. More specifically, we implement LAME with a RestNet-18 encoder with and without Hi-Vec for the CIFAR-10-C dataset with 20% outliers, which depicts the standard setting of our ablation studies. We find that with the implementation of Hi-Vec, the standard implementation of LAME can be improved from 68.4% to 79.7% accuracy, which is significantly higher than the source baseline of 57.3%. This highlights the ability of Hi-Vec to also integrate with parameter-free methods apart from backpropagation-based methods and motivates us to further investigate this in future work.
A.11 Algorithms
We provide the details about source model training in Algorithm 2. The detailed test-time algorithm is in the main paper.
@@ B.2 Results across common test-time adaptation setting
Table 6: Comparisons on common test-time adaptation setting. for CIFAR10-C and CIFAR100-C. Integrating Hi-Vec improves the common methods and achives the the best results (bold)
@@ B.3 Inference time
During source training, Hi-Vec introduces a small parameter overhead of approximately 5% compared to standard implementations. For example, for ResNet-18, the encoder has approximately 11M parameters while the added layers contribute approximately 0.5M parameters. This overhead scales modestly to larger models, even when implemented for intermediate layers such as recent Matformer [18]. At test-time, Hi-Vec needs to compute multiple gradient norms for layer selection. However, to adapt to target data, it doesn’t introduce any additional backward passes or parameters needed for backpropagation. Specifically, Hi-Vec uses only one of the hierarchical linear layers for backpropagation and during inference, ensuring no additional parameters are introduced. In Table 7, we compare the inference times for common test-time adaptation methods with and without Hi-Vec integration based on a single Nvidia A100 GPU. Hi-Vec introduces a modest increase in compute time for Tent and SAR, with Tent + Hi-Vec taking 1m 25s compared to 31s for Tent and SAR + Hi-Vec taking 1m 42s compared to 50s for SAR. For STAMP, Hi-Vec results in a negligible change. Overall, Hi-Vec achieves robust adaptation with minimal computational overhead, making it useful for test-time adaptation.
@@ C.1 Additional datasets information
We train source models respectively on the source datasets as per the training and evaluation procedure as in  [82, 47, 96, 26]. Akin to these common methods, the source datasets to train the respective source models include CIFAR-10, CIFAR-100, Waterbirds, and ImageNet as datasets. We utilize ResNet-18 and ResNet-50 with batch norm as the encoders. Our experimental setup involves multiple datasets at test time to evaluate the proposed method under two distinct types of diverse distribution shifts at test time. First, we consider outlier-aware scenarios where each test batch contains samples from Cifar-10-c [34], Cifar-100-c [34], or ImageNet-C [34] datasets combined with a significant proportion of outlier datasets, including LSUN-C [96], SVHN-C [96], TinyImageNet-C [96], Places-365-C [96], and Textures-C [96]. For the in-distribution data, Cifar-10-c [34] and Cifar-100-c [34] include 15 corruption types (e.g., Gaussian noise, motion blur, fog) applied at five severity levels, with Cifar-10-c [34] containing 10,000 images per corruption and Cifar-100-c [34] containing 10,000 images per class. ImageNet-C [34] serves as a large-scale benchmark with similar corruption types across 1,000 categories. We follow [96] to generate the outlier datasets. The outlier datasets are derived by applying the same 15 corruption types at the highest severity level of 5 to datasets such as LSUN, SVHN, TinyImageNet, Places-365, and Textures. In this setup, a single test batch contains both in-distribution samples from the corrupted CIFAR or ImageNet datasets and out-of-distribution samples from these corrupted outlier datasets. Next, we evaluate robustness to spurious correlations using datasets specifically designed to introduce misleading or non-causal correlations. The Waterbirds [72] dataset consists of 11,788 images where spurious correlations arise due to the background (e.g., water or land) being associated with specific bird species. Similarly, ColoredMNIST [72] modifies the MNIST dataset by introducing color as a spurious feature correlated with digit labels, consisting of 70,000 images across 10-digit classes. We train source models on Cifar-10, Cifar-100, Waterbirds, and ImageNet, respectively, using ResNet-18 and ResNet-50 as encoders. For ablations and additional experiments, we integrate Hi-Vec with re-implemented baselines from their official GitHub repositories.
@@ C.2 Hyperparameters for the baselines
C.3 Results for the main paper with standard deviations.
We also provide the standard deviations for the main results from the main paper.
Noise SVHN-C LSUN-C TinyImageNet-C
Methods Acc AUC H-score Acc AUC H-score Acc AUC H-score ACC AUC H-score
MRL [46] 59.7 65.7 62.6 59.7 64.4 62.0 59.7 61.9 60.2 59.7 64.1 61.8
BN Stats [60] 72.9 68.6 70.6 78.7 75.3 76.9 79.4 79.4 79.4 79.0 72.9 75.8
EATA [61] 72.9 68.5 70.6 78.8 75.3 76.9 79.4 79.4 79.4 78.9 73.1 75.9
CoTTA [84] 77.3 62.4 67.3 81.6 78.6 80.1 82.2 84.2 83.2 81.9 75.3 78.4
RoTTA [97] 77.6 74.3 75.6 78.4 76.0 77.2 78.8 79.5 79.1 78.6 73.3 75.8
SoTTA [25] 77.8 51.7 61.6 79.3 72.8 75.9 79.8 77.9 78.8 79.6 72.6 75.9
OWTTT [51] 62.3 64.4 58.5 66.1 75.3 69.6 63.1 78.9 68.5 56.3 58.8 56.2
Tent [82] 77.4 48.7 59.7 80.8 54.9 65.1 81.2 62.3 70.4 81.1 65.6 72.4
SAR [62] 72.9 68.5 70.6 78.7 75.3 76.9 79.4 79.4 79.4 79.0 72.9 75.8
STAMP  [96] 77.9 83.2 80.1 82.3 79.2 80.6 83.5 86.3 84.8 82.6 74.9 78.5
Tent + Hi-Vec 80.7 $\pm$0.2 $\blacktriangle$ 63.0 $\pm$0.2 $\blacktriangle$ 70.5 $\pm$0.2 $\blacktriangle$ 81.7 $\pm$0.1 $\blacktriangle$ 55.4 $\pm$0.1 $\blacktriangle$ 66.0 $\pm$0.1 $\blacktriangle$ 81.7 $\pm$0.2 $\blacktriangle$ 62.8 $\pm$0.2 $\blacktriangle$ 71.0 $\pm$0.2 $\blacktriangle$ 82.5 $\pm$0.1 $\blacktriangle$ 65.8 $\pm$0.1 $\blacktriangle$ 73.2 $\pm$0.1 $\blacktriangle$
SAR + Hi-Vec 77.7 $\pm$0.1 $\blacktriangle$ 63.8 $\pm$0.1 $\blacktriangle$ 70.7 $\pm$0.1 $\blacktriangle$ 82.5 $\pm$0.1 $\blacktriangle$ 73.3 $\pm$0.1 $\blacktriangle$ 77.7 $\pm$0.1 $\blacktriangle$ 80.2 $\pm$0.2 $\blacktriangle$ 79.9 $\pm$0.2 $\blacktriangle$ 80.0 $\pm$0.2 $\blacktriangle$ 83.0 $\pm$0.1 $\blacktriangle$ 69.7 $\pm$0.1 $\blacktriangle$ 75.7 $\pm$0.1 $\blacktriangle$
STAMP + Hi-Vec 83.6 $\pm$0.1 $\blacktriangle$ 91.4 $\pm$0.1 $\blacktriangle$ 87.3 $\pm$0.1 $\blacktriangle$ 85.7 $\pm$0.1 $\blacktriangle$ 82.7 $\pm$0.1 $\blacktriangle$ 84.2 $\pm$0.1 $\blacktriangle$ 84.3 $\pm$0.2 $\blacktriangle$ 86.9 $\pm$0.2 $\blacktriangle$ 85.5 $\pm$0.2 $\blacktriangle$ 86.5 $\pm$0.1 $\blacktriangle$ 81.1 $\pm$0.1 $\blacktriangle$ 83.7 $\pm$0.1 $\blacktriangle$
MRL [46] 41.4 60.3 47.8 41.4 61.0 47.3 41.4 58.0 48.3 41.4 63.3 48.4
BN Stats [60] 45.8 80.9 58.4 52.7 72.5 60.9 53.7 73.8 62.0 53.2 68.6 59.7
EATA [61] 55.2 86.1 67.1 58.1 75.6 65.6 58.8 77.2 66.7 58.6 70.7 64.0
CoTTA [84] 47.0 83.4 59.9 53.7 73.2 61.8 54.3 76.9 63.6 54.5 68.1 60.4
RoTTA [97] 47.9 54.0 49.4 47.3 67.0 55.3 48.3 69.5 56.7 47.8 65.5 55.0
SoTTA [25] 54.4 53.3 52.8 53.6 70.3 60.7 54.4 70.8 61.4 53.9 68.4 60.1
OWTTT [51] 47.1 70.3 56.2 53.9 74.3 62.3 54.5 73.5 62.5 54.2 68.5 60.4
Tent [82] 47.9 55.8 51.2 54.4 70.4 61.2 55.4 72.4 62.7 55.0 68.6 60.9
SAR [62] 57.5 88.6 68.9 59.2 65.2 61.9 60.5 73.5 66.3 60.8 72.1 65.9
STAMP  [96] 57.9 98.4 72.8 63.7 82.1 71.7 63.7 82.6 71.9 63.9 75.5 69.2
Tent + Hi-Vec 54.9 $\pm$0.2 $\blacktriangle$ 68.2 $\pm$0.2 $\blacktriangle$ 60.1 $\pm$0.2 $\blacktriangle$ 54.7 $\pm$0.2 $\blacktriangle$ 73.9 $\pm$0.2 $\blacktriangle$ 62.3 $\pm$0.2 $\blacktriangle$ 57.1 $\pm$0.2 $\blacktriangle$ 72.7 $\pm$0.2 $\blacktriangle$ 63.9 $\pm$0.2 $\blacktriangle$ 55.3 $\pm$0.1 $\blacktriangle$ 69.9 $\pm$0.1 $\blacktriangle$ 61.1 $\pm$0.1 $\blacktriangle$
SAR + Hi-Vec 57.9 $\pm$0.1 $\blacktriangle$ 89.2 $\pm$0.1 $\blacktriangle$ 69.4 $\pm$0.1 $\blacktriangle$ 54.9 $\pm$0.1 $\blacktriangle$ 73.4 $\pm$0.1 $\blacktriangle$ 62.8 $\pm$0.1 $\blacktriangle$ 60.9 $\pm$0.2 $\blacktriangle$ 73.9 $\pm$0.2 $\blacktriangle$ 66.7 $\pm$0.2 $\blacktriangle$ 62.1 $\pm$0.1 $\blacktriangle$ 74.4 $\pm$0.1 $\blacktriangle$ 68.4 $\pm$0.1 $\blacktriangle$
STAMP + Hi-Vec 58.2 $\pm$0.1 $\blacktriangle$ 89.6 $\pm$0.1 $\blacktriangle$ 73.5 $\pm$0.1 $\blacktriangle$ 64.4 $\pm$0.1 $\blacktriangle$ 82.5 $\pm$0.1 $\blacktriangle$ 72.1 $\pm$0.1 $\blacktriangle$ 63.8 $\pm$0.2 $\blacktriangle$ 82.5 $\pm$0.2 $\blacktriangle$ 72.0 $\pm$0.2 $\blacktriangle$ 64.6 $\pm$0.1 $\blacktriangle$ 75.8 $\pm$0.1 $\blacktriangle$ 70.4 $\pm$0.1 $\blacktriangle$
Table 8: Results on adaptation with outlier datasets using Cifar-10-C and Cifar-100-C as target datasets with four outlier datasets. We report the baselines and use evaluation metrics as provided by Yu et al. [96] with ResNet-18. Our results are averaged over fine runs. Hi-Vec consistently improves performance (indicated by $\blacktriangle$  ) over the common methods and is the top-performer (bold)
Dataset Methods Acc (%) Worst-Group Acc (%)
ColoredMNIST Source 63.40 20.05
MRL [46] 85.24 60.31
Tent [82] 57.06 9.80
MEMO [100] 63.77 6.23
SENTRY [69] 63.23 15.78
EATA [61] 60.81 17.98
SAR [63] 58.37 12.36
DeYO [47] 78.24 67.39
SAR + Hi-Vec 62.71 $\pm$0.3 $\blacktriangle$ 15.68 $\pm$0.3 $\blacktriangle$
DeYO + Hi-Vec 79.53 $\pm$0.2 $\blacktriangle$ 68.62 $\pm$0.2 $\blacktriangle$
WaterBirds Source 83.16 64.90
MRL [46] 85.24 60.31
Tent [82] 82.95 54.14
MEMO [100] 82.34 50.47
SENTRY [69] 85.77 60.90
EATA [61] 82.38 52.38
SAR [63] 82.60 53.41
DeYO [47] 87.42 73.92
SAR + Hi-Vec 83.25 $\pm$0.3 $\blacktriangle$ 55.61 $\pm$0.3 $\blacktriangle$
DeYO + Hi-Vec 89.53 $\pm$0.2 $\blacktriangle$ 77.23 $\pm$0.2 $\blacktriangle$
Table 9: Results for spurious correlation datasets using ColoredMNIST and WaterBirds with ResNet-18. We report baselines provided by [47]. Our results are averaged over five runs. Hi-Vec improves ($\blacktriangle$  ) the common methods and performs the best (bold).
Places365-C Textures-C
Methods ACC AUC H-score ACC AUC H-score
Source 18.2 61.6 26.1 18.2 54.6 25.8
MRL [46] 18.4 61.9 28.3 18.4 54.6 27.4
BN Stats [60] 31.1 67.7 41.1 31.6 61.2 40.7
EATA [61] 46.4 72.6 56.0 46.4 62.2 52.8
CoTTA [84] 33.8 66.9 43.5 34.2 60.7 42.8
RoTTA [97] 36.6 68.6 46.5 37.0 65.3 46.5
SoTTA [25] 41.7 67.8 50.7 41.8 60.3 48.8
OWTTT [51] 9.1 54.0 13.9 9.4 59.4 14.6
Tent [82] 34.9 51.8 39.5 39.0 48.6 42.0
SAR [62] 44.9 73.3 55.0 45.6 67.0 54.0
STAMP 46.4 77.7 57.6 46.5 71.9 56.2
Tent + Hi-Vec 35.4 $\pm$0.4 $\blacktriangle$ 52.0 $\pm$0.4 $\blacktriangle$ 42.1 $\pm$0.4 $\blacktriangle$ 39.4 $\pm$0.3 $\blacktriangle$ 59.1 $\pm$0.3 $\blacktriangle$ 47.2 $\pm$0.3 $\blacktriangle$
SAR + Hi-Vec 45.3 $\pm$0.2 $\blacktriangle$ 73.8 $\pm$0.2 $\blacktriangle$ 56.1 $\pm$0.2 $\blacktriangle$ 46.2 $\pm$0.3 $\blacktriangle$ 67.4 $\pm$0.3 $\blacktriangle$ 54.8 $\pm$0.3 $\blacktriangle$
STAMP + Hi-Vec 46.9 $\pm$0.3 $\blacktriangle$ 77.9 $\pm$0.3 $\blacktriangle$ 58.5 $\pm$0.3 $\blacktriangle$ 46.8 $\pm$0.2 $\blacktriangle$ 71.7 $\pm$0.2 $\blacktriangle$ 56.6 $\pm$0.2 $\blacktriangle$
Table 10: Results on adaptation with outlier datasets using Imagenet-C with Places365-C and Textures-C as outlier datasets and ResNet-50. We report baselines and use evaluation metrics as provided by [96]. Our results are averaged over five runs. The conclusion is similar, improvement ($\blacktriangle$  ) over common methods and performs the best (bold).
@@ Appendix D Additional Related work
In addition to the aforementioned applications, test-time adaptation has also been used in vision language models [4, 102, 78, 1], continual learning [37, 57, 95], classification that consider practical scenarios [27, 12, 15, 42, 80, 49, 20]. Moreover, recently, [41] proposed a framework that uses linear mode connectivity for adapted models during inference, with standard ResNet models.